Test the following numbers for divisibility by 6,9 and 11( Do not divide or factories)

a) 6 184 046

b)21 589 434


1
Expert's answer
2021-05-10T13:45:10-0400

a) 6184046 is the even number.

Sum of digits: 6+1+8+4+0+4+6=29 is not divisible by 3.

Thus, 6184046 is not divisible by 6.

Sum of odd digits: 6+8+0+6=20.

Sum of even digits: 1+4+4=9.

Since sum of odd digits does not equal to sum of even digits, 6184046 is not divisible by 11.


b) 21589434 is the even number.

Sum of digits: 2+1+5+8+9+4+3+4=36 is divisible by 3.

Thus, 21589434 is divisible by 6.

Sum of odd digits: 2+5+9+3=19.

Sum of even digits: 1+8+4+4=17.

Since sum of odd digits does not equal to sum of even digits, 21589434 is not divisible by 11.